# Curve fitting on a quantum annealer for an advanced navigation method

### v2

@@ III.3 Chebyshev polynomials
Besides the triangular functions, we shall also consider Chebyshev polynomials of the first kind as defined in Eq. 3, $\phi_{j}=T_{j}$, i.e., the fit function is now given by
@@
In Tab. 4, we show our results for fitting a quadratic Chebyshev polynomial (i.e., $m=3$) to the quadratic sample data 22; again, $n=64$, $d=8$, and $p=d-1=7$, which results in a QUBO matrix of size $24\times 24$. While the tabu solution is in acceptable agreement with the classical one, the quantum annealing solution deviates significantly. We find the same result for the cubic data (also with varying numerical coefficients) with a cubic Chebyshev polynomial ($m=4$) as the fit function. In Fig. 3, we show the RMSE of the tabu (red squares) and quantum annealing (green diamonds) solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$, respectively, as a function of $d$, the number of digits that is used for the binary fixed-point format of the expansion coefficients. Furthermore, we show the RMSE of the classical solution, which is independent of $d$, as a gray dashed line. The underlying sample data is the cubic one and the Chebyshev fit is of degree three, i.e., $m=4$. We plot against $d$ because we would like to keep the degree of the Chebyshev polynomial fixed while letting the QUBO matrix increase in size. While the tabu solution shows the expected behavior (cf. Fig. 1), we are not able to obtain a satisfying quantum annealing solution for problem sizes larger than $20\times 20$. This is in stark contrast to our findings for piecewise-linear curve fitting using triangular functions.
@@
This seems to hint toward the limitations of D-Wave’s current quantum annealing hardware for curve fitting with polynomials formulated as a QUBO. More precisely, the corresponding QUBO problem, which can be seen as a graph with logical qubits representing nodes that are connected by edges with weights given by the matrix elements of the QUBO matrix, needs to be mapped onto the so-called working graph of the quantum annealing device. Therefore, one might need several physical qubits to represent one logical qubit. Particularly for highly-connected (dense) QUBO problems, this could yield an Ising model that exhibits a complicated low-energy landscape where the ground state is difficult to access. This would explain why we were not successful in obtaining a feasible Chebyshev approximation as soon as its QUBO matrix is larger than the above mentioned size of about $20\times 20$. Such a matrix is, as apparent from the right diagram of Fig. 4, indeed relatively highly connected.
The amount of nonvanishing off-diagonal elements of the QUBO matrix for least-squares curve fitting is solely determined by the standardized functions $\phi_{j}$ (keeping the underlying to-be-approximated data fixed); see Eq. 6. More precisely, in case of the Chebyshev approximation we deal with polynomials $T_{j}$ whose support are the whole real line,
@@
for all $j$. Therefore, the resulting QUBO problem corresponds to a rather highly connected graph, and the matrix has many nonzero elements even far off the diagonal; see the right diagram of Fig. 4 for a graphical depiction. On the other hand, the QUBO matrix for a piecewise-linear approximation is always a band matrix, see the left diagram of Fig. 4, because the triangular functions $\Lambda_{j}$ have compact support,
$\operatorname{supp}(\Lambda_{j})=\begin{cases}[\tilde{x}_{0},\tilde{x}_{1}]&\text{if\leavevmode\nobreak\ }j=0\,,\\[2.5pt] [\tilde{x}_{m-2},\tilde{x}_{m-1}]&\text{if\leavevmode\nobreak\ }j=m-1\,,\\[2.5pt] [\tilde{x}_{j-1},\tilde{x}_{j+1}]&\text{otherwise}\,.\end{cases}$ (32)
@@
    the Chebyshev solution is less accurate compared to the piecewise-linear approximation for feasible cases of the size of the QUBO problem.
@@ IV.1 General formalism
More formally speaking, we consider a deterministic Markov decision process that starts at time $t=0$ and terminates at $t=T$, where the time evolution is discrete in terms of integer time steps: $t=0,\dotsc,T$. The process is characterized by sets $X_{t}\subset\mathbb{R}^{k}$ and $U_{t}\subset\mathbb{R}^{l}$, where $k,l\in\mathbb{N}$, that contain all states the process can be in and all actions that can be performed in order to change the current state, respectively, at time step $t$ [^4] In general, the actions depend on the state $x_{t}\in X_{t}$ of the process, i.e., $U_{t}=U_{t}(x_{t})$. However, in terms of notation, we omit this dependency for the the sake of brevity.
@@
where $D\colon X_{T}\to\mathbb{R}$ encodes terminal costs that occur at the end of the process. The value function is a measure for how cost-effectively the decision process has been if started at time step $t$ with initial state $x_{t}$ and running until the end by virtue of a given sequence of actions $u_{t},\dotsc,u_{T-1}$. The states $x_{t+1},\dotsc,x_{T}$ in Eq. (34) are generated according to the temporal evolution defined by the process’ dynamics 33.
@@
Solving the recursively-coupled set of equations is a nontrivial task. Fortunately, the Bellman equation suggests a practical algorithm to obtain the globally optimal solution. Typically, one is interested in the optimal control $u_{0}^{\ast},\dotsc,u_{T-1}^{\ast}$ of the whole process. The solution strategy consists of a backward (value) and forward (policy) iteration; see, e.g., Ref. [6]. First, one computes the function values of the optimal value function starting from the last time step and then moving sequentially backward in time, i.e., for all $t=T,\dotsc,0$ and for all states $x_{t}\in X_{t}$, the values $V_{t}^{\ast}(x_{t})$ are computed according to Eq. 36. Second, the optimal actions are obtained via
@@ IV.2 Toward an optimized speed profile
In principle, the Bellman equation as given in Eq. 36 can be used right away to compute an optimized speed profile. The mapping onto that task is, for example, accomplished by the following setup:
@@
  • $C_{t}(x_{t},u_{t})$ are the fuel costs that arise if the vessel moves with speed $u_{t}$ while being at position $x_{t}$ at time $t$.
@@ IV.3 Proof of principle
for all $t$ [^5] From a physics point of view, $u_{t}$ and $w$ are velocities. Thus, the transition function actually reads $x_{t+1}=x_{t}+(u_{t}-w(x_{t}))\hskip 0.83344pt\Delta t$ with $\Delta t$ denoting the time difference between two time steps. Since we are working solely with dimensionless quantities and integer time steps, $\Delta t=1$ and it is omitted for the sake of brevity. Here, $w\colon X_{t}\to\mathbb{R}$ denotes the magnitude of the projected sea current with respect to the direction of travel at position $x_{t}$ on the trajectory, $v_{\textup{max}}$ represents the maximal possible speed through water, and $\alpha\in\mathbb{R}_{>0}$ is a numerical parameter to fine-tune the terminal costs.
Upon inserting these definitions into Eq. 34 and writing the final state as $x_{T}=x_{s}+\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))$ as a sum of actions starting at time $s$ in state $x_{s}$, which follows directly from Eq. 40, the value function for the whole process reads
@@
$\phantom{=\;}+\alpha\biggl{[}\ell-\frac{1}{\ell}\biggl{(}x_{s}-\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))\biggr{)}\biggr{]}^{2}$
$\phantom{=\;}+1\,.$ (42)
Without loss of generality (and for numerical convenience), we assume that the sets of states and actions $X_{t}$ and $U_{t}$, respectively, are the same for every time step, i.e., independent of $t$. It is then straightforward to find the optimal actions, which are the roots of $\nabla_{u}V_{s}$ (for fixed $x_{s}$). Regarding the gradient of the value function, we find
@@
Now, we shall solve this optimization problem by numerically solving the Bellman equation (36) and subsequently using Eq. 37 to find the optimal actions. For the sake of simplicity, we set $w(x_{t})=0$ for all $x_{t}\in X_{t}$ and all $t$, and consider a voyage with $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. Neglecting phases of (de)acceleration and without currents, the optimal solution is obviously $u_{0}^{\ast}=\dotso=u_{3}^{\ast}=25$, i.e., moving (trivially) with constant velocity. Approaching this problem in terms of the value function 41, we find from Eq. (43) the analytical solution
@@
In the course of that, the traditional/naive approach discretizes the state space. Hence, after the value iteration, the resulting optimal value function $V_{t}^{\ast}$ is defined only on that discrete state space grid. Consequently, this poses constraints on the possible discretizations of the action space for the policy iteration 37 because for a given $x_{t}$, $u_{t}$ must be chosen such that $x_{t+1}=\Gamma_{t}(x_{t},u_{t})$ lies on the state space grid, too, or the off-grid result $\Gamma_{t}(x_{t},u_{t})$ has to be “snapped” onto the nearest grid point. Both approaches induce a discretization error that manifests in the final (pseudo-)optimal policy. The in that way obtained optimal policy most likely differs from the true, globally-optimal solution—it is only optimal with respect to the discretization of the state space. In our case, this would result in a pseudo-optimal speed profile different from the analytical solution 44.
@@
with $m\in\mathbb{N}$, and $\phi_{j}\colon X_{t}\to\mathbb{R}$ denote suitable standardized (basis) functions. Here, we used the circumflex notation to indicate that the optimal value function is approximated, and the superscript of the expansion coefficients emphasizes that there is a separate approximation for each time step. For the sake of brevity, we omit the dependence of $\hat{V}_{t}^{\ast}$ on the coefficient vector $c$ for the remainder of this section. The value iteration takes the same form as before (see Section IV.1) but with the subtle yet important difference that the optimal value function is approximated at every time step. Starting from the end of the horizon and moving backward in time, the approximation at time step $t$ is then used for $t-1$. We start at the finite horizon $t=T$ with the boundary condition $\hat{V}_{T}^{\ast}(x_{T})=D(x_{T})$ and then proceed backward with
@@
for $t=T-1,\dotsc,0$. Here, it is worth to emphasize that at every time step, this value iteration uses only a subset of the state space due to the involved approximation, which allows for an efficient treatment of large state spaces. In the literature, this idea is also known as fitted value iteration [10]. Eventually, the resulting sequence $\hat{V}_{T}^{\ast},\dotsc,\hat{V}_{0}^{\ast}$ of optimal approximated value functions is used in the usual policy iteration 37. Since these optimal value functions are now available for (in principle) arbitrary state space variables thanks to the continuous (yet fitted) nature of Eq. 45, the restrictions for the discretization of the action space are greatly relaxed.
@@
Now, we would like to present our numerical results. We solve the Bellman equation with its ingredients given in Eqs. (38)–(41). Again, we neglect currents and set $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. The analytical solution according to Eq. 44 is therefore given by
$u_{0}^{\ast}=\dotso=u_{3}^{\ast}=\frac{25}{1+1/\alpha}\,,$ (47)
which is the optimal solution for an arbitrary $\alpha<\infty$. The Bellman equation is solved grid-based as well as with approximating the optimal value function in the sense of a fitted value iteration. For the latter, the approximation task—which is nothing but curve fitting—is accomplished classically as well as using the QUBO formulation of Section II.2 solved on a quantum annealer. [^6] Again, the Advantage2 system but with 50 runs, 256 reads per run, and an annealing time of 1000 microseconds. We employ triangular functions for the expansion 45, $\phi_{j}=\Lambda_{j}$ (see Eq. 4), with $m=9$, $d=9$, and $p=8$ for the QUBO.
Figure 5: Visualization of the value function e approximation using different solution strategies (upper row) and their corresponding mean squared errors (MSE) (lower row). The left column shows the results for time slice $t=3$, while the right column presents the same analysis for the first time slice $t=0$. The upper panels depict the approximated value functions, whereas the lower panels show the associated approximation errors. Here and in the following figures, we use the MSE in contrast to the RMSE as defined in Eq. 26.
  • Analytic: Approaching the problem in terms of the value function 41, we find from $\nabla_{u}V_{s}=0$ (see Eq. 43) the analytical solution.
In Tab. 5, we display our results for the optimal policies obtained by means of the different solution strategies. We have an example where the grid-based solution is indeed less optimal than the results that employ the fitted value iteration with the approximation solved classically by matrix inversion and QUBO-based via a tabu search. This is caused, as explained earlier, by the discretization and forcing off-grid values onto on-grid ones. Moreover, using a QUBO solved on a quantum annealer for the curve fitting in course of the fitted value iteration yields a worse optimal policy.
@@
Figure 6: MSE of the reconstructed solution $\hat{V}^{\ast}$ compared to the reference solution $V^{\ast}$. They carry no time index because we averaged over all time slices. Left: Convergence behavior of the MSE with respect to the number of discretization points $n$ of the state space for different regularization parameters $\alpha\in\{10,100,1000\}$. Right: Error decay with respect to the number of basis functions $m$ for fixed discretization $n=50$ and varying $\alpha$. The horizontal dashed-dotted line indicates the target error threshold, where the approximation solution with respect to the number of basis functions $m$ starts to be more accurate than the grid solution.
@@
Figure 7: Error comparison of the two solution approaches, quantum annealing and tabu search, at different precision levels ($d=7$ and $d=9$), as a function of the number of basis functions $m$. The left panel shows the results for the first time slice in the fitted value iteration process, while the right panel displays the average performance over all time slices.
@@
It should be noted that the choice of $m$ and $d$ results in an $81$-dimensional QUBO problem. The annealer does not solve this problem size accurately enough, which is consistent with the findings in Section III. Fig. 5 shows that the solution of the quantum annealing already exhibits a significant error at the very first time slice, preventing the method from reaching the optimal solution. Due to error propagation, this issue is further exacerbated in the subsequent time slices. The reasons lie in the size of the search space and the specific energy landscape formed by the binary representation of the coefficients.
@@
The example calculated above does not provide a clear picture of the behavior if one varies the number of triangular functions $m$. Therefore, we varied the latter and analyzed the resulting errors. As shown in Fig. 6, increasing the number of basis functions leads to improved approximation results. This holds also for other values of $\alpha$. At this point, all calculations shown in Fig. 6 are classical solutions of the QUBO formulation of the approximation task for $\hat{V}_{t}^{\ast}$ using Gaussian elimination. Basically, this procedure corresponds to computing a matrix inverse. The resulting solution can be interpreted as the best possible outcome within the span of the chosen basis. In a following figure, we further investigate how different solvers approximate this best solution. In summary, increasing the number of basis functions effectively reduces the approximation error. The remaining challenge is to ensure that the solvers are capable of achieving this improved level of solution quality.
@@
Our analysis shows in Fig. 7 that as the number of basis functions—and thus the problem size—increases, the solutions generally deviate further from the best possible solution, referred to as the inverse solution (indicated in blue). This trend holds for the tabu search but is especially pronounced in the case of quantum annealing. Starting from six basis functions, the problem size becomes large enough for the approximation error to increase significantly. For smaller problem sizes, the quantum annealing yields approximation results that are comparable to those obtained with the classical solution or the tabu search. However, beyond a certain problem size, this similarity breaks down as clearly illustrated in Fig. 7.

### v3

@@ III.3 Chebyshev polynomials
Besides the triangular functions, we shall also consider Chebyshev polynomials of the first kind as defined in Eq.˜3, $\phi_{j}=T_{j}$, i.e., the fit function is now given by
@@
Table 4: Optimal expansion coefficients of the individual Chebyshev polynomials of a quadratic fit, obtained classically via matrix inversion and by solving a QUBO via tabu search as well as quantum annealing (labeled “qa”). Also shown are the APE of the optimal coefficients obtained from the QUBO with respect to the classical ones.
Classical $0.544598$ 0.031799-0.031799 $0.499257$
QUBO (tabu) $0.523438$ 0.039063-0.039063 $0.484375$ $3.89$ $22.84$ $2.98$
QUBO (qa) $0.445313$ 0.101563-0.101563 $0.421875$ $18.23$ $219.39$ $15.50$
Figure 3: RMSE of the tabu search (red squares, $f_{\textup{tb}}^{\ast}$) and quantum annealing (green diamonds, $f_{\textup{qa}}^{\ast}$) solutions as a function of $d$. The RMSE of the exact, $d$-independent classical solution $f_{\textup{cl}}^{\ast}$ is shown as a gray dashed line. Lines between data points are shown to guide the eye. The underlying sample data set is the cubic one (Eq.˜24), and Chebyshev polynomials with $m=4$ are used for the approximation.
@@
In Tab.˜4, we show our results for fitting a quadratic Chebyshev polynomial (i.e., $m=3$) to the quadratic sample data 23; again, $n=64$, $d=8$, and p=d1=7p=d-1=7, which results in a QUBO matrix of size $24\times 24$. While the tabu solution is in acceptable agreement with the classical one, the quantum annealing solution deviates significantly. We find the same result for the cubic data (also with varying numerical coefficients) with a cubic Chebyshev polynomial ($m=4$) as the fit function. In Fig.˜3, we show the RMSE of the tabu (red squares) and quantum annealing (green diamonds) solutions $f_{\textup{tb}}^{\ast}$ and $f_{\textup{qa}}^{\ast}$, respectively, as a function of $d$, the number of digits that is used for the binary fixed-point format of the expansion coefficients. Furthermore, we show the RMSE of the classical solution, which is independent of $d$, as a gray dashed line. The underlying sample data is the cubic one and the Chebyshev fit is of degree three, i.e., $m=4$. We plot against $d$ because we would like to keep the degree of the Chebyshev polynomial fixed while letting the QUBO matrix increase in size. While the tabu solution shows the expected behavior (cf. Fig.˜1), we are not able to obtain a satisfying quantum annealing solution for problem sizes larger than $20\times 20$. This is in stark contrast to our findings for piecewise-linear curve fitting using triangular functions.
This seems to hint toward the limitations of D-Wave’s current quantum annealing hardware for curve fitting with polynomials formulated as a QUBO. More precisely, the corresponding QUBO problem, which can be seen as a graph with logical qubits representing nodes that are connected by edges with weights given by the matrix elements of the QUBO matrix, needs to be mapped onto the so-called working graph of the quantum annealing device. Therefore, one might need several physical qubits to represent one logical qubit. Particularly for highly-connected (dense) QUBO problems, this could yield an Ising model that exhibits a complicated low-energy landscape where the ground state is difficult to access. This would explain why we were not successful in obtaining a feasible Chebyshev approximation as soon as its QUBO matrix is larger than the above mentioned size of about $20\times 20$. Such a matrix is, as apparent from the right diagram of Fig.˜4, indeed relatively highly connected.
The amount of nonvanishing off-diagonal elements of the QUBO matrix for least-squares curve fitting is solely determined by the standardized functions $\phi_{j}$ (keeping the underlying to-be-approximated data fixed); see Eq.˜6. More precisely, in case of the Chebyshev approximation we deal with polynomials $T_{j}$ whose support are the whole real line,
@@
for all $j$. Therefore, the resulting QUBO problem corresponds to a rather highly connected graph, and the matrix has many nonzero elements even far off the diagonal; see the right diagram of Fig.˜4 for a graphical depiction. On the other hand, the QUBO matrix for a piecewise-linear approximation is always a band matrix, see the left diagram of Fig.˜4, because the triangular functions $\Lambda_{j}$ have compact support,
$\operatorname{supp}(\Lambda_{j})=\begin{cases}[\tilde{x}_{0},\tilde{x}_{1}]&\text{if\penalty 10000\ }j=0\,,\\[2.5pt] [\tilde{x}_{m-2},\tilde{x}_{m-1}]&\text{if\penalty 10000\ }j=m-1\,,\\[2.5pt] [\tilde{x}_{j-1},\tilde{x}_{j+1}]&\text{otherwise}\,.\end{cases}$ (33)
@@
    the Chebyshev solution is less accurate compared to piecewise-linear approximation for feasible cases of the size of the QUBO problem.
@@ IV.1 General formalism
More formally speaking, we consider a deterministic Markov decision process that starts at time $t=0$ and terminates at $t=T$, where the time evolution is discrete in terms of integer time steps: $t=0,\dotsc,T$. The process is characterized by sets $X_{t}\subset\mathbb{R}^{k}$ and $U_{t}\subset\mathbb{R}^{l}$, where $k,l\in\mathbb{N}$, that contain all states the process can be in and all actions that can be performed in order to change the current state, respectively, at time step tt. [^4] In general, the actions depend on the state $x_{t}\in X_{t}$ of the process, i.e., $U_{t}=U_{t}(x_{t})$. However, in terms of notation, we omit this dependency for the the sake of brevity.
@@
where $D\colon X_{T}\to\mathbb{R}$ encodes terminal costs that occur at the end of the process. The value function is a measure for how cost-effectively the decision process has been if started at time step $t$ with initial state $x_{t}$ and running until the end by virtue of a given sequence of actions $u_{t},\dotsc,u_{T-1}$. The states $x_{t+1},\dotsc,x_{T}$ in Eq.˜35 are generated according to the temporal evolution defined by the process’ dynamics 34.
@@
Solving the recursively-coupled set of equations is a nontrivial task. Fortunately, the Bellman equation suggests a practical algorithm to obtain the globally optimal solution. Typically, one is interested in the optimal control $u_{0}^{\ast},\dotsc,u_{T-1}^{\ast}$ of the whole process. The solution strategy consists of a backward (value) and forward (policy) iteration; see, e.g., Ref. [6]. First, one computes the function values of the optimal value function starting from the last time step and then moving sequentially backward in time, i.e., for all $t=T,\dotsc,0$ and for all states $x_{t}\in X_{t}$, the values $V_{t}^{\ast}(x_{t})$ are computed according to Eq.˜37. Second, the optimal actions are obtained via
@@ IV.2 Toward an optimized speed profile
In principle, the Bellman equation as given in Eq.˜37 can be used right away to compute an optimized speed profile. The mapping onto that task is, for example, accomplished by the following setup:
@@
  • $C_{t}(x_{t},u_{t})$ are the fuel costs if the vessel moves with speed $u_{t}$ while being at position $x_{t}$ at time $t$.
@@ IV.3 Proof of principle
for all tt. [^5] From a physics point of view, $u_{t}$ and $w$ are velocities. Thus, the transition function actually reads $x_{t+1}=x_{t}+(u_{t}-w(x_{t}))\hskip 0.83344pt\Delta t$ with $\Delta t$ denoting the time difference between two time steps. Since we are working solely with dimensionless quantities and integer time steps, $\Delta t=1$ and it is omitted for the sake of brevity. Here, $w\colon X_{t}\to\mathbb{R}$ denotes the magnitude of the projected sea current with respect to the direction of travel at position $x_{t}$ on the trajectory, $v_{\textup{max}}$ represents the maximal possible speed through water, and $\alpha\in\mathbb{R}_{>0}$ is a numerical parameter to fine-tune the terminal costs.
Upon inserting these definitions into Eq.˜35 and writing the final state as $x_{T}=x_{s}+\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))$ as a sum of actions starting at time $s$ in state $x_{s}$, which follows directly from Eq.˜41, the value function for the whole process reads
@@
$+\alpha\biggl[\ell-\frac{1}{\ell}\biggl(x_{s}-\sum_{t=s}^{T-1}(u_{t}-w(x_{t}))\biggr)\biggr]^{2}$
$+1\,.$ (43)
Without loss of generality (and for numerical convenience), we assume that the sets of states and actions $X_{t}$ and $U_{t}$, respectively, are the same for every time step, i.e., independent of $t$. It is then straightforward to find the optimal actions, which are the roots of $\nabla_{\hskip-0.68295ptu}V_{s}$ (for fixed $x_{s}$). Regarding the gradient of the value function, we find
@@
Now, we shall solve this optimization problem by numerically solving the Bellman equation 37 and subsequently using Eq.˜38 to find the optimal actions. For the sake of simplicity, we set $w(x_{t})=0$ for all $x_{t}\in X_{t}$ and all $t$, and consider a voyage with $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. Neglecting phases of (de)acceleration and without currents, the optimal solution is obviously $u_{0}^{\ast}=\dotso=u_{3}^{\ast}=25$, i.e., moving (trivially) with constant velocity. Approaching this problem in terms of the value function 42, we find from Eq.˜44 the analytical solution
@@
In the course of that, the traditional/naive approach discretizes the state space. Hence, after the value iteration, the resulting optimal value function $V_{t}^{\ast}$ is defined only on that discrete state space grid. Consequently, this poses constraints on the possible discretizations of the action space for the policy iteration 38 because for a given $x_{t}$, $u_{t}$ must be chosen such that $x_{t+1}=\Gamma_{t}(x_{t},u_{t})$ lies on the state space grid, too, or the off-grid result $\Gamma_{t}(x_{t},u_{t})$ has to be “snapped” onto the nearest grid point. Both approaches induce a discretization error that manifests in the final (pseudo-)optimal policy. The in that way obtained optimal policy most likely differs from the true, globally-optimal solution—it is only optimal with respect to the discretization of the state space. In our case, this would result in a pseudo-optimal speed profile different from the analytical solution 45.
@@
with $m\in\mathbb{N}$, and $\phi_{j}\colon X_{t}\to\mathbb{R}$ denote suitable standardized (basis) functions. Here, we used the circumflex notation to indicate that the optimal value function is approximated, and the superscript of the expansion coefficients emphasizes that there is a separate approximation for each time step. For the sake of brevity, we omit the dependence of $\hat{V}_{t}^{\ast}$ on the coefficient vector $c$ for the remainder of this section. The value iteration takes the same form as before (see Section˜IV.1) but with the subtle yet important difference that the optimal value function is approximated at every time step. Starting from the end of the horizon and moving backward in time, the approximation at time step $t$ is then used for t1t-1. We start at the finite horizon $t=T$ with the boundary condition $\hat{V}_{T}^{\ast}(x_{T})=D(x_{T})$ and then proceed backward with
@@
for $t=T-1,\dotsc,0$. Here, it is worth to empasize that at every time step, this value iteration uses only a subset of the state space due to the involved approximation, which allows for an efficient treatment of large state spaces. In the literature, this idea is also known as fitted value iteration [10]. Eventually, the resulting sequence $\hat{V}_{T}^{\ast},\dotsc,\hat{V}_{0}^{\ast}$ of optimal approximated value functions is used in the usual policy iteration 38. Since these optimal value functions are now available for (in principle) arbitrary state space variables thanks to the continuous (yet fitted) nature of Eq.˜46, the restrictions for the discretization of the action space are greatly relaxed.
@@
Now, we would like to present our numercial results. We solve the Bellman equation with its ingredients given in Eqs. (39)–(42). Again, we neglect currents and set $\ell=100$, $v_{\textup{max}}=50$, and $T=4$. The analytical solution according to Eq.˜45 is therefore given by $u_{0}^{\ast}=\dotso=u_{3}^{\ast}=25\hskip 0.83344pt\alpha/(\alpha+1)$, which is the optimal solution for an arbitrary $\alpha<\infty$. The Bellman equation is solved grid-based as well as with approximating the optimal value function in the sense of a fitted value iteration. For the latter, the approximation task—which is in this case nothing but curve fitting—is accomplished classically as well as using the QUBO formulation of Section˜II.2 solved on a quantum annealer. [^6] Again, the Advantage2 QPU but with 50 runs, 256 reads per run, and an annealing time of 1000 microseconds. We employ triangular functions for the expansion 46, $\phi_{j}=\Lambda_{j}$ (see Eq.˜4), with $m=9$, $d=9$, and p=d1=8p=d-1=8 for the QUBO.
@@
Figure 5: Visualization of the value function approximation using different solution strategies (upper row) and their corresponding mean squared errors (MSE) (lower row). The left column shows the results for time slice $t=3$, while the right column presents the same analysis for the first time slice $t=0$. The upper panels depict the approximated value functions, whereas the lower panels show the associated approximation errors. Here and in the following figures, we use the MSE in contrast to the RMSE as defined in Eq.˜27.
  • Analytic: Approaching the problem in terms of the value function 42, we find the analytical solution from $\nabla_{\hskip-0.68295ptu}V_{s}=0$.
@@
Figure 6: MSE of the reconstructed solution $\hat{V}^{\ast}$ compared to the reference solution $V^{\ast}$. They carry no time index because we averaged over all time slices. Left: Convergence behavior of the MSE with respect to the number of discretization points $n$ of the state space for different regularization parameters $\alpha\in\{10,100,1000\}$. Right: Error decay with respect to the number of basis functions $m$ for fixed discretization $n=50$ and varying $\alpha$. The horizontal dash-dotted line indicates the target error threshold, where the approximation solution with respect to the number of basis functions $m$ starts to be more accurate than the grid solution.
@@
Figure 7: Error comparison of the two solution approaches, quantum annealing and tabu search, at different precision levels ($d=7$ and $d=9$), as a function of the number of basis functions. The left panel shows the results for the first time slice in the fitted value iteration process, while the right panel displays the average performance over all time slices.
In Tab.˜5, we display our results for the optimal policies obtain by means of the different solution strategies. We have an example where the grid-based solution is indeed less optimal than the results that employ the fitted value iteration with the approximation solved classically by matrix inversion and QUBO-based via a tabu search. This is caused, as explained earlier, by the discretization and forcing off-grid values onto on-grid ones. Moreover, using a QUBO solved on a quantum annealer for the curve fitting in course of the fitted value iteration yields a worse optimal policy.
@@
It should be noted that the choice of $m$ and $d$ results in an $81$-dimensional QUBO problem. The annealer does not solve this problem size accurately enough, which is consistent with our findings in Section˜III. Figure 5 shows that the solution of the quantum annealing already exhibits a significant error at the very first time slice, preventing the method from reaching the optimal solution. Due to error propagation, this issue is further exacerbated in the subsequent time slices. The reasons lie in the size of the search space and the specific energy landscape formed by the binary representation of the coefficients.
@@
The example calculated above does not provide a clear picture of the behavior if one varies the number of triangular functions $m$. Therefore, we varied the latter and analyzed the resulting errors. As shown in Fig.˜7, increasing the number of basis functions leads to improved approximation results. This holds also for other values of $\alpha$. At this point, all calculations shown in Fig.˜7 are classical solutions of the QUBO formulation of the approximation task for $\hat{V}_{t}^{\ast}$ using Gaussian elimination. Basically, this procedure corresponds to computing a matrix inverse. The resulting solution can be interpreted as the best possible outcome within the span of the chosen basis. In a following figure, we further investigate how different solvers approximate this best solution. In summary, increasing the number of basis functions effectively reduces the approximation error. The remaining challenge is to ensure that the solvers are capable of achieving this improved level of solution quality.
@@
Our analysis shows in Fig.˜7 that as the number of basis functions—and thus the problem size—increases, the solutions generally deviate further from the best possible solution, referred to as the inverse solution (indicated in red). This trend holds for the tabu search but is especially pronounced in the case of quantum annealing. Starting from six basis functions, the problem size becomes large enough for the approximation error to increase significantly. For smaller problem sizes, the quantum annealing yields approximation results that are comparable to those obtained with the classical solution or the tabu search. However, beyond a certain problem size, this similarity breaks down as clearly illustrated in Fig.˜7.